Ivanhoe Monograms sells stadium blankets that have been monogrammed with high school and university emblems. The blankets retail for $45 throughout the country to loyal alumni of over 3,500 schools. Ivanhoe’s variable costs are 41% of sales; fixed costs are $118,000 per month.

Assume that variable costs increase to 45% of the current sales price and fixed costs increase by $13,400 per month. If Ivanhoe were to raise its sales price 11% to cover these new costs, but the number of blankets sold were to drop by 5%, what would be the new annual operating income?
Answer #1

Selling price = $45 + 11% = $49.95

Variable cost = $45 X 45% = $20.25

Fixed cost = $118,000 + $13,400 = $131,400

Blankets sold = 3,500 - 5% = 3,325

Sales (3,325 X $49.95) $166,084
(Less): Variable cost (3,325 X $20.25) $67,331
Contribution margin $98,753
(Less): Fixed cost $131,400
Net operating loss ($32,647)